# Invisible Languages of the LLM Universe

### v1

@@ 5.3 Limitations and Future Directions
Methodological limitations: First, our digitality measurement relies on publicly accessible web data, potentially missing private communication in messaging applications. If Invisible Giant speakers use their languages extensively via WhatsApp, SMS, or encrypted platforms, our analysis underestimates actual digital use. However, from an LLM training perspective, private communication does not contribute to model development, so our measurement captures the dimension relevant for AI representation.
Second, the four-category classification uses median splits, creating discrete boundaries where gradients exist. Languages near category boundaries may be misclassified, and the framework does not capture within-category variation – not all Strongholds enjoy equal digital presence, nor do all Invisible Giants face identical challenges. Future work should develop continuous measures of the representation gap alongside categorical classifications.
Third, we lack longitudinal data on how vitality-digitality gaps evolve. Are gaps widening or narrowing? Do digital interventions successfully narrow gaps for targeted languages? Cross-sectional analysis cannot answer these causal questions. Longitudinal tracking of language trajectories across the vitality-digitality space would enable causal inference about what drives languages from one quadrant to another.
Theoretical extensions: Future research should examine intersections of linguistic marginalization with other forms of oppression. How does digital language inequality interact with disability (especially for signed languages’ digital representation)? With gender, particularly for languages where speaker populations are gendered? With age, as youth language innovations occur primarily in digital spaces? These intersectional analyses remain underdeveloped.
Platform-specific language ideologies warrant deeper investigation. Different platforms encode different linguistic norms – Twitter’s character limits advantage logographic systems, TikTok’s audio-visual primacy benefits languages with media presence, LinkedIn’s professional register expectations advantage languages with established business communication. How do platform affordances systematically advantage or disadvantage specific languages? We need granular sociotechnical analysis of how architectural choices embed linguistic assumptions.
Community-centered research methodologies remain rare. Our study, like most, analyzes patterns from external positions without deep engagement with affected language communities in research design or interpretation. Developing genuine co-production methodologies where speakers shape research questions, interpret findings, and control data governance represents a methodological frontier essential for decolonizing research practice.
Empirical extensions: The framework’s generalizability across domains requires testing. We developed it for languages; Saurabh Khanna’s broader Invisible Information project applies similar logic to armed conflicts and scholarly research. Does the vitality-digitality logic extend to other knowledge domains – traditional medicine, indigenous land management practices, oral historical archives? Systematic testing across domains would validate or refine theoretical claims about digital filtering mechanisms.
@@
Regional deep dives would complement global analysis. Why do specific Invisible Giants remain underrepresented despite demographic strength? Detailed case studies examining language-specific barriers – orthographic debates, institutional politics, diaspora dynamics – would illuminate mechanisms obscured in aggregate analysis. Combining computational measurement with ethnographic depth would strengthen explanatory power.
@@ 5.4 Practical Implications
For AI Developers: Current LLM training treats language representation as a data availability problem – scrape more web pages, digitize more books. Our analysis reveals this approach is insufficient. The 2,000 Invisible Giants have speakers, they have knowledge to contribute, but they lack digital infrastructure and institutional support that would enable content creation at scales sufficient for LLM training.
@@
Targeted investment in digital infrastructure for Invisible Giants is required: orthography standardization where needed, keyboard interfaces, spell-checkers, text-to-speech systems. These are prerequisites for content creation, not merely nice-to-have additions. Community-controlled data trusts enabling speakers to govern how language data is collected, used, and monetized would shift power dynamics from extractive to collaborative.
Evaluation metrics must center non-English linguistic features – tone marking accuracy, morphological complexity handling, pragmatic particle usage – rather than English-centric benchmarks like BLEU scores that penalize linguistic differences as errors. Performance equity should be measured not just by aggregate metrics but by whether models serve speakers of Invisible Giants as effectively as they serve English speakers.
@@
For Policymakers: Language planning must integrate digital dimensions. Granting a language official status without digital infrastructure investment produces hollow recognition. Policies should mandate that government digital services support languages spoken by significant populations, fund localization of open-source software and educational platforms, and establish accountability mechanisms requiring AI companies to report representation gaps and remediation efforts.
Digital Language Rights should extend existing linguistic human rights frameworks into digital domains. Just as the Universal Declaration of Linguistic Rights (1996) establishes rights to use one’s language in education and public life, contemporary frameworks must recognize rights to digital language resources, AI services in one’s language, and participation in knowledge production digitized by platforms.
International funding mechanisms should prioritize Invisible Giants over Fading Voices for digital investment. While all endangered languages merit documentation, the 2,000 Invisible Giants offer highest social return on investment – millions of speakers exist who could immediately benefit from digital resources. This is not abandoning Fading Voices but strategically allocating limited resources where impact potential is greatest.
For Researchers: Methodological decolonization requires fundamental shifts in how we conduct research. Co-designing studies with speaker communities rather than extracting data, compensating language consultants at professional rates, sharing model weights and tools with communities providing training data, publishing in open-access venues with translations into studied languages – these practices should become norms, not exceptions.
@@
Evaluating research impact by community-defined outcomes (language vitality metrics, economic benefits, cultural continuity) rather than only academic citations would better align incentives with social benefit. Research excellence should incorporate community endorsement as a criterion, recognizing that scholarship about marginalized communities should serve those communities’ self-determined goals.
Funding agencies could accelerate these shifts by requiring invisibility impact assessments for digitization grants – explicitly addressing how proposed work will or will not narrow representation gaps – and prioritizing proposals co-designed with affected language communities over researcher-initiated projects.
@@ 5.5 Toward Linguistic Justice in AI
The Invisible Giants paradox – millions speaking, none listening digitally – crystallizes the crisis. Solutions require confronting uncomfortable truths: English dominance in AI is not natural or technologically necessary but reflects historical power structures that continue shaping which languages receive investment, recognition, and computational support.
Optimistically, concentrated investment could rapidly narrow gaps. The methodology exists: measure representation gaps, prioritize Invisible Giants, invest in community-controlled infrastructure, develop multilingual models with performance parity guarantees. The economics are feasible – comprehensive digital infrastructure for 2,000 Invisible Giants would cost a fraction of single LLM training runs measuring in hundreds of millions of dollars.
Pessimistically, current incentives point opposite directions. Commercial AI development optimizes for wealthy markets speaking Stronghold languages. Academic research chases state-of-the-art performance on English benchmarks, with multilingual work often treated as secondary. Platform business models reward engagement metrics that favor content in dominant languages. Without structural intervention – regulatory requirements, funding conditionalities, community organizing – market forces will accelerate inequality.
The choice before us is clear: will AI systems democratize access to information, education, and economic opportunity across linguistic boundaries? Or will they accelerate language shift, compressing human diversity toward English-centered homogeneity? The answer depends on whether we treat the vitality-digitality gap as a technical resource allocation problem or as the digital-epistemic injustice it truly represents – requiring not just more data, but fundamental redistribution of power, resources, and recognition.
@@ 6 Conclusion
We have demonstrated that linguistic inequality in Large Language Models is not incidental but structural, reflecting continuities from colonial-era hierarchies to contemporary AI development. Of the world’s 7,613 languages, roughly 2,000 Invisible Giants with millions of speakers remain digitally underrepresented, concentrating in formerly colonized regions at rates nine times higher than Europe.
This pattern constitutes digital-epistemic injustice: the systematic exclusion of marginalized linguistic communities from AI-mediated knowledge production through denial of digital infrastructure and epistemic authority. As LLMs become infrastructure for education, commerce, and governance, this exclusion threatens to accelerate language shift, compressing human linguistic diversity toward homogenous English-centered multilingualism.
Yet this outcome is not inevitable. Concentrated investment in digital infrastructure for systematically under-resourced languages, community-controlled data governance, and decolonial AI development practices could narrow gaps within a generation. The question is one of political will: do we accept AI systems that reproduce colonial hierarchies, or do we demand technology that honors and amplifies the full richness of human linguistic diversity?
The invisible languages of the LLM universe need not remain invisible. Making them visible requires recognizing that their exclusion was never accidental – and that inclusion requires not just more data, but fundamentally reimagining who AI serves and whose knowledge counts.
@@ Acknowledgments
We acknowledge research funding from the Fair and Resilient Societies Working Group and the Amsterdam School of Communication Research at the University of Amsterdam. We also acknowledge the countless language communities whose knowledge and labor make AI systems possible, yet who rarely benefit from the technologies built on their linguistic resources.

### v2

@@ 5.2 Limitations and Future Directions
Methodological limitations
First, our digitality measurement relies on publicly accessible web data, potentially missing private communication in messaging applications. If Invisible Giant speakers use their languages extensively via WhatsApp, SMS, or encrypted platforms, our analysis underestimates actual digital use. However, from an LLM training perspective, private communication does not contribute to model development, so our measurement captures the dimension relevant for AI representation. Second, the four-category classification uses median splits, creating discrete boundaries where gradients exist. Languages near category boundaries may be misclassified, and the framework does not capture within-category variation – not all Strongholds enjoy equal digital presence, nor do all Invisible Giants face identical challenges. Future work should develop continuous measures of the representation gap alongside categorical classifications. Third, we lack longitudinal data on how vitality-digitality gaps evolve. Are gaps widening or narrowing? Do digital interventions successfully narrow gaps for targeted languages? Cross-sectional analysis cannot answer these causal questions. Longitudinal tracking of language trajectories across the vitality-digitality space would enable causal inference about what drives languages from one quadrant to another.
Theoretical extensions
Future research should examine intersections of linguistic marginalization with other forms of oppression. How does digital language inequality interact with special needs (especially for signed languages’ digital representation)? With gender, particularly for languages where speaker populations are gendered? With age, as youth language innovations occur primarily in digital spaces? These intersectional analyses remain underdeveloped. Platform-specific language ideologies warrant deeper investigation. Different platforms encode different linguistic norms – Twitter’s character limits advantage logographic systems, TikTok’s audio-visual primacy benefits languages with media presence, LinkedIn’s professional register expectations advantage languages with established business communication. How do platform affordances systematically advantage or disadvantage specific languages? We need granular socio-technical analysis of how architectural choices embed linguistic assumptions. Community-centered research methodologies remain rare. Our study, like most, analyzes patterns from external positions without deep engagement with affected language communities in research design or interpretation. Developing genuine co-production methodologies where speakers shape research questions, interpret findings, and control data governance represents a methodological frontier essential for decolonizing research practice.
Empirical extensions
@@
The framework’s generalizability across domains requires testing, since we only developed it for languages, while similar frameworks can extend towards assessing low visibility of certain armed conflicts and scholarly research. Does the vitality-digitality logic extend to other knowledge domains – traditional medicine, land management practices, oral historical archives? Systematic testing across domains would validate or refine theoretical claims about digital filtering mechanisms. Regional deep dives would complement global analysis. Why do specific Invisible Giants remain underrepresented despite demographic strength? Detailed case studies examining language-specific barriers – orthographic debates, institutional politics, diaspora dynamics – would illuminate mechanisms obscured in aggregate analysis. Combining computational measurement with ethnographic depth would strengthen explanatory power here.
@@ 5.3 Practical Implications
For AI Developers
@@
Current LLM training treats language representation as a data availability problem – scrape more web pages, digitize more books. Our analysis reveals this approach is insufficient. The 2,000 Invisible Giants have speakers, they have knowledge to contribute, but they lack digital infrastructure and institutional support that would enable content creation at scales sufficient for LLM training. Targeted investment in digital infrastructure for Invisible Giants is required: orthography standardization where needed, keyboard interfaces, spell-checkers, text-to-speech systems. These are prerequisites for content creation, not merely nice-to-have additions. Community controlled data trusts enabling speakers to govern how language data is collected, used, and monetized would shift power dynamics from extractive to collaborative. Evaluation metrics must center non-English linguistic features – tone marking accuracy, morphological complexity handling, pragmatic particle usage – rather than English-centric benchmarks like BLEU scores that penalize linguistic differences as errors. Performance equity should be measured not just by aggregate metrics but by whether models serve speakers of Invisible Giants as effectively as they serve English speakers.
For Policymakers
@@
Language planning must integrate digital dimensions. Granting a language official status without digital infrastructure investment produces hollow recognition. Policies should mandate that government digital services support languages spoken by significant populations, fund localization of open-source software and educational platforms, and establish accountability mechanisms requiring AI companies to report representation gaps and remediation efforts. Digital Language Rights should extend existing linguistic human rights frameworks into digital domains. Just as the Universal Declaration of Linguistic Rights (1996) establishes rights to use one’s language in education and public life, contemporary frameworks must recognize rights to digital language resources, AI services in one’s language, and participation in knowledge production digitized by platforms. International funding mechanisms should prioritize Invisible Giants for digital investment. These offer highest social return on investment – millions of speakers flourish who could immediately benefit from digital resources.
For Researchers
@@
Methodological decolonization requires fundamental shifts in how we conduct research. Co-designing studies with speaker communities rather than extracting data, compensating language consultants at professional rates, sharing model weights and tools with communities providing training data, publishing in open-access venues with translations into studied languages – these practices should become norms, not exceptions. Evaluating research impact by community-defined outcomes (language vitality metrics, economic benefits, cultural continuity) rather than only academic citations would better align incentives with social benefit. Research excellence should incorporate community endorsement as a criterion, recognizing that scholarship about marginalized communities should serve those communities’ self-determined goals. Funding agencies could accelerate these shifts by requiring visibility impact assessments for digitization grants – explicitly addressing how proposed work will or will not narrow representation gaps – and prioritizing proposals co-designed with affected language communities over top-down ones.
@@ 5.4 Toward Linguistic Justice in AI
The Invisible Giants paradox, populations numbering in the millions, whose speakers are effectively unheard in digital spaces – encapsulates the core of the current crisis. Addressing this problem requires acknowledging a set of uncomfortable empirical and historical facts: the predominance of English in contemporary AI systems is neither an inherent linguistic inevitability nor a technical necessity. Rather, it is a manifestation of enduring geopolitical and socioeconomic power asymmetries that continue to determine which languages receive sustained investment, institutional recognition, and robust computational support.
From an optimistic standpoint, targeted and adequately resourced interventions could substantially reduce these disparities. The requisite methodological toolkit largely already exists: systematically quantify linguistic representation gaps; prioritize structurally marginalized, high-vitality languages; invest in community-governed digital and computational infrastructures; and design multilingual models with explicit guarantees of performance parity or fairness across languages. From a realistic view though, prevailing incentive structures are misaligned with these goals. Commercial AI development is primarily optimized for high-income markets whose users predominantly speak globally dominant stronghold languages. Academic research agendas frequently prioritize incremental improvements on English-centric benchmarks, while multilingual and low-resource language work is often formulated as secondary or peripheral. Platform-based business models privilege engagement metrics that systematically favor content in already dominant languages. In the absence of deliberate structural interventions – such as regulatory mandates, conditional public funding mechanisms, and sustained community-based organizing – market dynamics are likely to further intensify linguistic and digital inequalities.
The resulting policy and design choice is stark: Will emerging AI technologies be deployed to democratize access to information, education, and economic opportunity across linguistic boundaries, or will they function as accelerants of language shift, compressing linguistic diversity into an increasingly English-centered monoculture? The outcome hinges on whether the gap between linguistic vitality and digital representation is conceptualized merely as a technical resource allocation problem, or as a form of digital and epistemic injustice. The latter framing underscores that closing this gap requires not only additional data and infrastructure, but also a more profound redistribution of power, resources, and epistemic recognition across linguistic communities.
@@ 6 Conclusion
We demonstrate that linguistic inequality in Large Language Models (LLMs) is not incidental but structurally embedded, reproducing historical continuities from colonial language hierarchies through to contemporary AI development. Of the world’s 7,613 attested languages, approximately 2,000 can be classified as Invisible Giants – languages with millions of speakers that remain substantially underrepresented in digital ecosystems. This underrepresentation is disproportionately concentrated in formerly colonized regions, where rates of linguistic exclusion are estimated to be orders of magnitude higher than in Europe.
We conceptualize this configuration as a form of digital-epistemic injustice: the systematic marginalization of specific linguistic communities within AI-mediated regimes of knowledge production, enacted through constrained access to digital infrastructure and the systematic devaluation of their epistemic contributions. As LLMs increasingly function as core socio-technical infrastructure for education, commerce, governance, and cultural production, such exclusion is likely to intensify ongoing processes of language shift. This dynamic promotes convergence toward a homogenized, English-dominant model of multilingualism and, in turn, accelerates the erosion of global linguistic diversity.
This trajectory, however, is not immutable. Targeted and sustained investment in digital infrastructure for structurally under-resourced languages, the institutionalization of community-governed data stewardship, and the adoption of explicitly decolonial principles in AI design and governance have the potential to significantly mitigate these disparities within a generational timescale. The central question is political rather than purely technical: whether societies will accept AI systems that reproduce and normalize colonial linguistic hierarchies, or instead demand technological infrastructures that recognize, sustain, and amplify the full spectrum of human linguistic diversity.
The invisible languages within the current LLM ecosystem need not remain invisible. Rendering them visible requires acknowledging that their exclusion has been historically and structurally produced rather than accidental, and recognizing that meaningful inclusion entails more than the quantitative expansion of training datasets. It necessitates a fundamental reconfiguration of whose interests AI systems are designed to serve and whose knowledge systems are accorded legitimacy and authority.
@@ Acknowledgments
We acknowledge research funding for this project from the Amsterdam School of Communication Research at the University of Amsterdam.